OpenStudy (republican31):

Factor 5x^2 -7x-6 I know something about how to factor, but I'm not sure what to do because the LC isn't 1.

OpenStudy (princeharryyy):

5x^2-10x +3x - 6 => 5x(x-2) +3(x-2) => (5x+3)(x-2) => x=2, -5/3

OpenStudy (republican31):

Oh ok thank you!!! I've wondered forever how to solve these equations with a leading coefficient that's not one. Thank you!

OpenStudy (princeharryyy):

sorry, it's-3/5. I wrote it wrong. correct it.

OpenStudy (princeharryyy):

when leading coefficient is not one, you have to multiply the first one to the third one. suppose it is..... ax^2 + bx +c you have to multiply a with c and make b as the sum of two numbers in such a form that the product of of both comes out to be a*b

OpenStudy (republican31):

And then factor by grouping, yes?

OpenStudy (princeharryyy):

see here 5*(-6) = -30 and -10+3 = -7 -10*3 = -30 which is equal to 5*(-6) = -30

OpenStudy (princeharryyy):

and that's it.

OpenStudy (princeharryyy):

i hope you got it @Republican31 did u?

OpenStudy (republican31):

Yep! Thank you!

OpenStudy (arnavguddu):

5x^2 - 7x -6 Product = coeff of sq term and constant term = 5 * -6 = -30 factors of 30 = 2 * 3* 5 factors of -30 = 2 * 3* 5 and odd number of these factors will be negative like -2, -3, 5, 2, -3, 5, etc now combine these factors such that their sum is -7 => -2*5 +3 = -7. here the factors used are (-2, 3, 5) now given eq = 5x^2 -2*5*x +3x -6 = 5x^2 - 10x + 3x - 6 = 5x(x-2) + 3(x-2) = (x-2) (5x-2) done!!!